• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 14:43:58 on October 13, 2009

    Entry number 294740

    keyword=Compton Rate Adjustment

    Around 11:30 this morning, the Compton rates started to increase. The
    MCC crew was contacted and Yan Wang made an adjustment to MBH2I03V to lower
    the Compton rate.

    See Elog # 1492073.

    Afterwards, the Program Deputy Yves Roblin said that by tweaking the above
    magnet, they can decrease the Compton rate increases, which is being
    caused by the drifts. This should not affect the other halls and an
    expert will keep an eye on this and adjust it when needed. The drifts
    appear to be on a 12 hour time scale.

    A copy of this log entry has been emailed to: camsonne,paschke,mfriend,dwang,dalton,mcnulty



    ELOG # 1492073